OpenStudy (anonymous):

CAlculus Help! find the derivative of the function. g(x)=3arccos(x/2)

OpenStudy (anonymous):

are u familiar with the derivative of \[\cos^{-1}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[g(x)=3cos^{-1}(\frac{x}{2})\] \[cos(\frac{g(x)}{3})=cos(cos^{-1}(\frac{x}{2})=\frac{x}{2}\] if you take the derivative of that, you can find your answer much easier using chain rule without having to memorize arccos derviatives.

OpenStudy (anonymous):

\[g'(x) = 3[-\frac{1/2}{\sqrt{1-(\frac{1}{2})^{2}}}]\] that took a while, but this is how i'm learning to do these derivatives

OpenStudy (anonymous):

but i HAVE to memorize the arccos derivatives!

OpenStudy (anonymous):

umm okay can u exaplin step by step? pleas?

OpenStudy (anonymous):

2le or me?

OpenStudy (anonymous):

i mean explain?

OpenStudy (anonymous):

well you take the derivative and plug in the the values to the formula.

OpenStudy (anonymous):

why did u take out the 3?

OpenStudy (anonymous):

the (3) is the constant multiple rule

OpenStudy (anonymous):

its a constant being multiplied so its carried out

OpenStudy (anonymous):

o okay. :]

OpenStudy (anonymous):

wait and the 1/2? y is it on top?

OpenStudy (anonymous):

so why is the 1/2 on top again?

zepdrix (zepdrix):